# Global approximations to the error function of real argument for vectorized computation

### v1

@@ Abstract.
2020 Mathematics Subject Classification:
Primary 33B20, 65D20; Secondary 33F05, 33-04
1. Introduction
The error function [6] of real argument
(1.1) $\operatorname{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}\exp\left(-x^{2}\right)\,\mathrm{d}x$
shows up in many mathematical models of physical and other phenomena (far too many even to be listed here), and its numerical evaluation can be a bottleneck of a computational simulation. Standard mathematical libraries of C and FORTRAN implement it since at least 2008, and use diverse approximations for some defined ranges of $x$, favoring accuracy over speed, but it can be helpful to have a faster though slightly less accurate implementaion. The vector instructions of modern processors promise a speedup of up to 16 times, but then a branch-free code is needed to harness them. In some physical models [1, 5] the error function divided by its argument, a well-behaved even function
(1.2) $F_{0}(x)=\frac{\operatorname{erf}(x)}{x}$
@@ 1. Introduction
We have found global closed-form approximations to both functions (1.1) and (1.2) in terms of addition, multiplication, division, and square root — with or without also using the exponential function — where the accuracy can be systematically improved by taking more polynomial terms with optimized coefficients, reaching 128 bits of precision and stopping there.
We confess having found our approximation formulas by general mathematical arguments using the natural intelligence of our own mind, but then we had to make sure this had not been done before. We see in the literature that the approximations to the error function have been developed since the early days of computation [9, 7, 8, 10, 2, 11], but ours still seems to be new. (We cannot review all such works here as it may grow into a study in the psychology of mathematics.)
2. Approximations
We begin with a transformation of the error function (1.1)
(2.1) $\operatorname{erf}(x)=\frac{x}{\sqrt{x^{2}+\phi\left(x^{2}\right)}}$
in terms of a new function $\phi(s)$, the $x$ in the numerator in (2.1) makes it ideal also for the function (1.2). Looking at its explicit form
(2.2) $\phi(s)=\frac{s}{\bigl{[}\operatorname{erf}\bigl{(}\sqrt{s}\bigr{)}\bigr{]}^{2}}-s,$
@@ 2. Approximations
(2.5) $\lim_{s\to\infty}\phi(s)=\frac{2}{\sqrt{\pi}}\sqrt{s}\exp(-s).$
@@
(2.7) $\psi_{N}(s)=\frac{\sum_{m=0}^{N+1}A_{mN}s^{m}}{1+\sum_{n=1}^{N}B_{nN}s^{n}}\approx\psi(s)$
can be made. The conditions (2.3), (2.4), and (2.5) now become
(2.8) $\psi(0)=\frac{\pi^{2}}{16},$
(2.9) $\psi^{\prime}(0)=\frac{5\pi^{2}}{24}-\frac{\pi}{2}=\frac{(5\pi-12)\pi}{24},$
(2.10) $\lim_{s\to\infty}\psi(s)=\frac{4}{\pi}s,$
and the rational function (2.7) can be easily constrained to satisfy them.
@@
(2.11) $\exp(-s)\approx\textstyle\left(1+\sum_{n=1}^{N}(2^{-K}s)^{n}b_{n}/n!\right)^{-2^{K}}$
@@
(2.12) $\phi_{MN}^{(K)}(s)=\left(\frac{\sum_{m=0}^{M}A_{mMN}^{(K)}s^{m}}{1+\sum_{n=1}^{N}B_{nMN}^{(K)}s^{n}}\right)^{2^{K}}\approx\phi(s)$
to work strikingly well for the right $K$, $M$, and $N$, even without satisfying (2.5).
The coefficients in (2.7) and (2.12) can be optimized to minimize the maximum
(2.13) $E=\max\limits_{0<x<\infty}\bigl{|}\varepsilon(x)\bigr{|}$
@@
(2.14) $\varepsilon(x)=\frac{f(x)}{\operatorname{erf}(x)}-1$
of the approximation $f(x)$ based on (2.1) with $\phi(s)$ either from (2.6) and (2.7) or from (2.12). In practice, this can be done by solving the system of equations
(2.15) $\left\{\begin{array}[]{ccrll}\varepsilon(x_{i})&=&-\varepsilon(x_{i+1}),&x_{i}<x_{i+1},&i=1,\dots,L,\\ \varepsilon^{\prime}(x_{i})&=&0,&&i=1,\dots,L+1,\end{array}\right.$
for $L$ variables: $L=2N-1$ for (2.7) with (2.8), (2.9), and (2.10); or $L=M+N-1$ for (2.12) with (2.3) and (2.4). The starting values of the parameters can be taken first from the minimization of the least-squares ($p=1$)
(2.16) $E^{(p)}=\int\limits_{0}^{\infty}\bigl{(}\varepsilon(x)\bigr{)}^{2p}\,\mathrm{d}x$
@@
3. Computations
We have written a computer code to determine the approximation coefficients using multiple-precision floating-point arithmetic, typically 256 bits. For the exponential-based approximation (2.7) we have found well-behaved solutions, with all coefficients positive, for up to $N=27$, but failed for $N=17,21,25$ where some $B_{nN}<0$. Table 1 shows the accuracy of these approximations given as $-\log_{2}E$, the number of significant bits, when the computation is done with a much higher bit precision, and we see an exponential convergence with $N$. For the exponential-free approximation (2.12) we do not claim to have worked through all combinations of $(M,N,K)$, nevertheless we have found 55 well-behaved solutions some of which are shown in Table 1 alongside the exponential-based solutions of comparable accuracy.
Table 1. Accuracy of approximations.
@@ 3. Computations
To study the effects of finite-precision arithmetic, and also as a way to share all our solutions, we have formatted the coefficients as C code (see supplementary material) to evaluate the approximations in 24-bit (mantissa) single, 53-bit double, 64-bit long double, and 113-bit quadruple precision, and to compare it to the standard library $\operatorname{erf}$ function. As our “standard” single- and double-precision approximations we have chosen those highlighted in Table 1, and the rounding errors add up to leave us with about 22, 21 (single) and 51, 48 (double) bits of precision.
To measure the computational speed, we have written C code (see supplementary material) for serial as well as 4-way double- and 8-way single-precision vectorized calculation, such that the GCC [4] compiler we use can translate it into either scalar or vector instructions. For the AVX2/FMA instruction set, we get a quite well-optimized machine code (see supplementary material) where the scalar and vector instructions nearly parallel each other, and run it on an AMD 3950X 16-core processor running at 3.5 GHz clock frequency with SMT turned off, 16 identical jobs in parallel to load all the cores. Timing the repeated evaluation of a function $f(x)$ over 512 equally-spaced values of $40\leq x<$ 4, for a total of about $2^{32}$ function calls, is used to estimate the number of processor clock cycles for one function value including load/store, call/return, and looping operations.
We compare the speed of the standard C library [3] implementation of $\exp$ and $\operatorname{erf}$ functions against the serial and vectorized code of our approximations. We also use this occasion to share our own vectorized single- and double-precision implementation of the exponential function where not the traditional Chebyshev but the direct uniform approximation to the $2^{x}$ function for $\tfrac{1}{2}\leq x\leq\tfrac{1}{2}$ is used.
Table 2. Measurements of computational speed.
@@
$\exp$ double glibc [3] 1 45
@@
$\operatorname{erf}$ double glibc [3] 1 83
$\operatorname{erf}$ double ours, Eq. 2.12 1 27 3.1
$\operatorname{erf}$ double ours, Eq. 2.12 4 34 3.2
$\operatorname{erf}$, $\exp$ double ours, Eq. 2.7 1 43
$\operatorname{erf}$, $\exp$ double ours, Eq. 2.7 4 65 2.6
$\exp$ single glibc [3] 1 19
@@
$\operatorname{erf}$ single glibc [3] 1 62
$\operatorname{erf}$ single ours, Eq. 2.12 1 16 3.9
$\operatorname{erf}$ single ours, Eq. 2.12 8 20 6.4
$\operatorname{erf}$, $\exp$ single ours, Eq. 2.7 1 26
$\operatorname{erf}$, $\exp$ single ours, Eq. 2.7 8 38 5.5

### v2

@@ Abstract
I Introduction
The error function Glaisher (1871) of real argument
$\operatorname{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}\exp\left(-x^{2}\right)\,\mathrm{d}x$ (1)
shows up in many mathematical models of physical and other phenomena (far too many even to be listed here), and its numerical evaluation can be a bottleneck of a computational simulation. Standard mathematical libraries of C and FORTRAN implement it since at least 2008, and use diverse approximations for some defined ranges of $x$, favoring accuracy over speed, but it can be helpful to have a faster though slightly less accurate implementaion. The vector instructions of modern processors promise a speedup of up to 16 times, but then a branch-free code is needed to harness them. In some physical models Boys (1950); Gill et al. (1996) the error function divided by its argument — a well-behaved even function
$F_{0}(x)=\frac{\operatorname{erf}(x)}{x}$ (2)
@@ I Introduction
We have found global closed-form approximations to both functions (1) and (2) in terms of addition, multiplication, division, and square root — with or without also using the exponential function — where the accuracy can be systematically improved by taking more polynomial terms with optimized coefficients, reaching 128 bits of precision and stopping there.
We confess having found our approximation formulas by general mathematical arguments using the natural intelligence of our own mind, but then we had to make sure this had not been done before. We see in the literature that the approximations to the error function have been developed since the early days of computation Hastings (1955); Hart (1957, 1966); Oldham (1968); Cody (1969); Shepherd and Laframboise (1981), but ours still seems to be new. (We cannot review all such works here as it may grow into a study in the psychology of mathematics.)
II Approximations
We begin with a transformation of the error function (1)
$\operatorname{erf}(x)=\frac{x}{\sqrt{x^{2}+\phi\left(x^{2}\right)}}$ (3)
in terms of a new function $\phi(s)$, the $x$ in the numerator in (3) makes it ideal also for the function (2). Looking at its explicit form
$\phi(s)=\frac{s}{\bigl[\operatorname{erf}\bigl(\sqrt{s}\bigr)\bigr]^{2}}-s,$ (4)
@@ II Approximations
$\lim_{s\to\infty}\phi(s)=\frac{2}{\sqrt{\pi}}\sqrt{s}\exp(-s).$ (7)
@@
$\psi_{N}(s)=\frac{\sum_{m=0}^{N+1}A_{mN}s^{m}}{1+\sum_{n=1}^{N}B_{nN}s^{n}}\approx\psi(s)$ (9)
can be made. The conditions (5), (6), and (7) now become
$\psi(0)=\frac{\pi^{2}}{16},$ (10)
$\psi^{\prime}(0)=\frac{(5\pi-12)\pi}{24},$ (11)
$\lim_{s\to\infty}\psi(s)=\frac{4}{\pi}s,$ (12)
and the rational function (9) can be easily constrained to satisfy them.
@@
$\exp(-s)\approx\textstyle\left(1+\sum_{n=1}^{N}(2^{-K}s)^{n}b_{n}/n!\right)^{-2^{K}}$ (13)
@@
$\phi_{MN}^{(K)}(s)=\left(\frac{\sum_{m=0}^{M}A_{mMN}^{(K)}s^{m}}{1+\sum_{n=1}^{N}B_{nMN}^{(K)}s^{n}}\right)^{2^{K}}\approx\phi(s)$ (14)
to work strikingly well for the right $K$, $M$, and $N$, even without satisfying (7).
The coefficients in (9) and (14) can be optimized to minimize the maximum
$E=\max\limits_{0<x<\infty}\bigl|\varepsilon(x)\bigr|$ (15)
@@
$\varepsilon(x)=\frac{f(x)}{\operatorname{erf}(x)}-1$ (16)
of the approximation $f(x)$ based on (3) with $\phi(s)$ either from (8) and (9) or from (14). In practice, this can be done by solving the system of equations
$\left\{\begin{array}[]{ccrll}\varepsilon(x_{i})&=&-\varepsilon(x_{i+1}),&x_{i}<x_{i+1},&i=1,\dots,L,\\ \varepsilon^{\prime}(x_{i})&=&0,&&i=1,\dots,L+1,\end{array}\right$. (17)
for $L$ variables: L=2N1L=2N-1 for (9) with (10), (11), and (12); or L=M+N1L=M+N-1 for (14) with (5) and (6). The starting values of the parameters can be taken first from the minimization of the least-squares ($p=1$)
$E^{(p)}=\int\limits_{0}^{\infty}\bigl(\varepsilon(x)\bigr)^{2p}\,\mathrm{d}x$ (18)
@@
III Computations
We have written a computer code to determine the approximation coefficients using multiple-precision floating-point arithmetic, typically 256 bits. For the exponential-based approximation (9) we have found well-behaved solutions, with all coefficients positive, for up to $N=27$, but failed for $N=17,21,25$ where some $B_{nN}<0$. Table 1 shows the accuracy of these approximations given as $\log_{2}E$, the number of significant bits, when the computation is done with a much higher bit precision, and we see an exponential convergence with $N$. For the exponential-free approximation (14) we do not claim to have worked through all combinations of $(M,N,K)$, nevertheless we have found 55 well-behaved solutions some of which are shown in Table 1 alongside the exponential-based solutions of comparable accuracy.
Table 1: Accuracy of approximations.
@@ III Computations
To study the effects of finite-precision arithmetic, and also as a way to share all our solutions, we have formatted the coefficients as C code Laikov (2015) to evaluate the approximations in 24-bit (mantissa) single, 53-bit double, 64-bit long double, and 113-bit quadruple precision, and to compare it to the standard library $\operatorname{erf}$ function. As our “standard” single- and double-precision approximations we have chosen those highlighted in Table 1, and the rounding errors add up to leave us with about 22, 21 (single) and 51, 48 (double) bits of precision.
To measure the computational speed, we have written C code Laikov (2015) for serial as well as 4-way double- and 8-way single-precision vectorized calculation, such that the GCC Free Software Foundation (2023) compiler we use can translate it into either scalar or vector instructions. For the AVX2/FMA instruction set, we get a quite well-optimized machine code Laikov (2015) where the scalar and vector instructions nearly parallel each other, and run it on an AMD 3950X 16-core processor clocked at 3.5 GHz with SMT turned off, 16 identical jobs in parallel to load all the cores. Timing the repeated evaluation of a function $f(x)$ over 512 equally-spaced values of $0\leq x<4$, for a total of about $2^{32}$ function calls, is used to estimate the number of processor clock cycles for one function value including load/store, call/return, and looping operations.
We compare the speed of the standard C library Free Software Foundation (2015) implementation of $\exp$ and $\operatorname{erf}$ functions against the serial and vectorized code of our approximations. We also use this occasion to share our own vectorized single- and double-precision implementation of the exponential function where not the traditional Chebyshev but the direct uniform approximation to the $2^{x}$ function for $\tfrac{1}{2}\leq x\leq\tfrac{1}{2}$ is used.
Table 2: Measurements of computational speed.
@@
$\exp$ double glibc Free Software Foundation (2015) 1 45
@@
$\operatorname{erf}$ double glibc Free Software Foundation (2015) 1 83
$\operatorname{erf}$ double ours, Eq. 14 1 27 3.1
$\operatorname{erf}$ double ours, Eq. 14 4 34 3.2
$\operatorname{erf}$, $\exp$ double ours, Eq. 9 1 43
$\operatorname{erf}$, $\exp$ double ours, Eq. 9 4 65 2.6
$\exp$ single glibc Free Software Foundation (2015) 1 19
@@
$\operatorname{erf}$ single glibc Free Software Foundation (2015) 1 62
$\operatorname{erf}$ single ours, Eq. 14 1 16 3.9
$\operatorname{erf}$ single ours, Eq. 14 8 20 6.4
$\operatorname{erf}$, $\exp$ single ours, Eq. 9 1 26
$\operatorname{erf}$, $\exp$ single ours, Eq. 9 8 38 5.5
@@ IV Conclusions
V Data availability
The approximation coefficients and computer codes are available online Laikov (2015).